### v2

# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

Abstract
Audio-aware large language models (ALLMs) have recently made great strides in understanding and processing audio inputs. These models are typically adapted from text-based large language models (LLMs) through additional training on audio-related tasks. However, this adaptation process presents two major limitations. First, ALLMs often suffer from catastrophic forgetting, where crucial textual capabilities like instruction-following are lost after training on audio data. In some cases, models may even hallucinate sounds that are not present in the input audio, raising concerns about reliability. Second, achieving cross-modal alignment between audio and language typically relies on large collections of task-specific question–answer pairs for instruction tuning, making it resource-intensive. To address these issues, previous works have leveraged the backbone LLMs to synthesize general-purpose, caption-style alignment data. In this paper, we propose a data generation framework that produces contrastive-like training data, designed to enhance ALLMs’ ability to differentiate between present and absent sounds. We further extend our approach to multi-audio scenarios, enabling the model to either explain differences between audio inputs or produce unified captions that describe all inputs, thereby enhancing audio-language alignment. We refer to the entire ALLM training framework as bootstrapping audio-language alignment via synthetic data generation from backbone LLMs (BALSa). Experimental results indicate that our method effectively mitigates audio hallucinations while reliably maintaining strong performance on audio understanding and reasoning benchmarks, as well as instruction-following skills. Moreover, incorporating multi-audio training further enhances the model’s comprehension and reasoning capabilities. Overall, BALSa offers an efficient and scalable approach to developing ALLMs.
Index Terms:
Audio-aware large language models, audio-language alignment
I Introduction
The development of large language models (LLMs) has garnered increasing attention from the research community. Building on this foundation, researchers have explored integrating different modalities into LLMs beyond their strong textual capabilities. For instance, incorporating visual or audio modalities enables these models to perform visual or auditory understanding, thereby extending their capabilities beyond text processing. This multimodal integration allows LLMs to leverage their inherent linguistic proficiency alongside their understanding of other modalities to perform multimodal reasoning and question answering.
In the audio-language domain, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] (ALLMs) refer to models that integrate audio processing capabilities into LLMs, enabling them to understand and generate responses based on auditory inputs. A critical step in developing these models is cross-modal alignment, where textual and auditory information are integrated to ensure effective reasoning across modalities.
To achieve such alignment in practice, the dominant approach relies on instruction tuning with large-scale audio-language datasets [1, 2, 3, 4, 5], typically composed of task-specific question-answer pairs. However, it is expensive to manually collect this kind of data, so some previous work [1, 2, 4, 5, 3] often turns to proprietary language models like ChatGPT [11] to synthesize it using metadata from existing audio corpora.
Beyond the data curation effort, another fundamental issue with existing ALLMs is catastrophic forgetting. Since most ALLMs are adapted from text-based LLMs, they inherently possess strong language understanding capabilities before being trained on audio tasks. However, prior research has reported that currently developed ALLMs often lose their textual capabilities after being trained on audio data. This includes a decline in instruction-following behavior [12], overfitting to pre-defined tasks [3], and limited generalization to unseen cross-modal reasoning tasks [3]. In some cases, the models may even mistakenly identify sounds that are not present in the audio [13].
To address both the challenge of catastrophic forgetting and the burden of data curation, there is a growing trend toward leveraging a backbone LLM to generate synthetic data for audio-language alignment [14, 15, 16, 17, 18, 19]. Here, the backbone LLM refers to the large language model that serves as the initialisation of the core component of ALLM architectures. Instead of relying on manually crafted or task-specific question-answer pairs, which are often generated using proprietary LLMs and complex rules, these approaches leverage the backbone LLM to synthesize training data in a prompt-driven, automatic, and extensible manner. The developers only need to design simple generation prompts, and the backbone LLM can flexibly generate diverse data with minimal human effort. While the conceptual direction of using backbone LLMs for data generation has been widely explored in the speech domain, this article is the first to focus on extending the idea to general audio tasks, demonstrating that it also offers advantages in simplicity, flexibility, and scalability in the audio domain.
This paper represents the framework, BALSa, short for Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs. Different from previous works using backbone LLMs for synthetic data generation, BALSa generates not only descriptive data, such as natural language descriptions of events occurring in the audio, but also contrastive data, which distinguish sounds that are present from those that are absent by simply varying the generation prompt [^1] The incorporation of the synthetic negative samples was first proposed in our previous conference paper, LISTEN (Learning to Identify Sounds Through Extended Negative Samples) [20]. This article extends that conference paper with a unified training framework, multi-audio modeling, and additional evaluations and experiments.. Like previous work of synthetic data generation, BALSa leverages the backbone LLM’s understanding of audio metadata, such as sound event tags, to synthesize training data. Consequently, we only need the prompts like Repeat the audio or List non-existent sounds in the audio to efficiently generate diverse audio-language data tailored to various learning objectives. BALSa offers a flexible and generalizable pipeline that facilitates audio-language alignment by learning from the behavioral patterns of the backbone LLM. This highlights BALSa’s potential as a general-purpose framework for synthesizing diverse alignment data for a wide range of audio understanding tasks. Building upon the flexibility of BALSa in generating diverse alignment data, our findings indicate that incorporating negative synthetic samples effectively mitigates hallucinations, thereby enhancing model reliability.
Furthermore, BALSa can be adapted and extended to various application scenarios. In this study, we introduce an extension of BALSa designed for multi-audio scenarios, where the backbone LLM independently generates explanations of the differences between multiple audio samples and a caption that describes both audio samples together, serving as training data. The objective is to expand the original alignment process into a twofold approach. First, a discrimination-based process that generates comparative explanations for differences between audio samples. Second, a description-based process that generates separate captions for each audio sample within the same response.
Our experimental results demonstrate that training the model on multi-audio data, whether through joint captioning or comparative discrimination, sufficiently enhances its overall performance. Interestingly, the difference between generating a single caption that encompasses both audio samples and producing comparative explanations for their differences is minimal. This suggests that simply learning to handle multiple audio samples at once is sufficient to improve the model’s comprehension, even without explicit comparative training. These findings indicate that exposing the model to multiple audio inputs, whether for captioning (descriptive learning) or discrimination (comparative learning), is a key factor in boosting performance. Furthermore, the BALSa-based approach leads to more efficient data usage. Our method achieves comparable or even better performance while using only 12% of the training data duration required by the previous state-of-the-art model [2].
To comprehensively evaluate model performance, we collect and construct multiple evaluation benchmarks, categorized into three major types: audio question answering, audio reasoning, and reliability and safety benchmarks. These benchmarks assess various aspects of the model’s capabilities, including basic audio understanding, advanced reasoning, and its reliability and safety in real-world applications. Through these evaluations, we compare our proposed approach against popular baseline ALLMs in terms of audio-language alignment. Moreover, we investigate catastrophic forgetting and hallucination issues to further assess model robustness and reliability. We provide illustrative examples on our demo page.  [^2] https://kuan2jiu99.github.io/Balsa
In conclusion, our contributions are outlined as follows:
  1. 1.
    We present BALSa (Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs). A critical feature of BALSa is a contrastive-style training strategy that incorporates synthetic negative samples, enabling models to distinguish clearly between present and absent sounds. Experimental results demonstrate that BALSa effectively mitigates audio hallucinations, thereby enhancing the model’s reliability in real-world applications.
  2. 2.
    Prior works [14, 15, 16, 17, 18, 19] have explored and validated the effectiveness of synthetic data generation using backbone LLMs. In these studies, the backbone LLM analyzes audio metadata and generates descriptions guided by designed prompts, but they focus primarily on the speech domain. This article is the first to apply this paradigm specifically to the non-speech audio domain. BALSa synthesizes audio-language alignment data using simple prompts from audio metadata, such as sound event tags, without requiring task-specific question-answer collections.
  3. 3.
    We extend BALSa to support multi-audio processing by introducing BALSa-MA (Multi-Audio), which applies BALSa to scenarios involving multiple audio samples. In this approach, the backbone LLM both compares audio samples by generating descriptions of their differences and provides separate captions for each sample within the same response. Experimental results indicate that exposing the model to multiple audio samples, whether through comparative explanations or joint captioning, enhances its audio understanding and improves its performance on evaluation benchmarks.
  4. 4.
    BALSa demonstrates high data efficiency, achieving performance comparable to or better than a widely adopted strong model [2], while using only 12% of the training data duration.
  5. 5.
    To comprehensively assess audio-aware large language models, we conduct experiments on existing benchmarks, reproduce missing evaluation sets, and introduce new test sets to explore previously unexamined aspects. Our evaluation covers three key areas: audio question answering, which assesses fundamental audio understanding; audio reasoning; and reliability and safety. Through these evaluations, we systematically compare BALSa with mainstream models, examining the effectiveness of audio-language alignment and addressing challenges like audio hallucinations and instruction adherence.
II Related Works
There is already extensive research on audio and speech-related large language models. Due to space constraints, this paper discusses only the studies directly relevant to our work. For a comprehensive overview of the broader framework of audio and speech LLMs, readers are encouraged to consult the survey papers [21, 22, 23, 24, 25, 26, 27, 28].
Large language models (LLMs) [29, 30, 31, 32, 33] can perform zero-shot learning through instruction tuning. This capability is expected to emerge when strong cross-modal alignment is achieved, particularly in scenarios where a text-based LLM is integrated with multi-modal encoders. Following this trend, audio-aware large language models [1, 2, 3, 4, 5, 6, 7, 8, 34] extend instruction tuning to audio-related tasks by combining audio encoders with language models and training them in a multi-stage process.
Qwen-Audio [1] adopts a two-stage training strategy. In the first stage, multi-task pre-training is conducted using a manually curated dataset that includes tasks such as acoustic scene classification, sound event classification, and audio question answering. To unify the training format across tasks, Qwen-Audio introduces task-specific tags that indicate the nature of each task. This stage aims to build a foundational understanding of audio signals. In the second stage, Qwen-Audio performs supervised instruction tuning to align the model more closely with human intent. To create training data, proprietary LLMs [11, 35, 36] are used to generate question–answer pairs based on text metadata, while additional audio-dialogue datasets are constructed using human annotations. Through this process, Qwen-Audio enhances the model’s ability in both audio understanding and reasoning. Qwen2-Audio [2] follows a similar approach, focusing on collecting high-quality supervised fine-tuning data through meticulous curation. SALMONN [3] also employs the same two-stage pipeline. During instruction tuning, it prioritizes tasks based on their relevance and manually collects task-specific datasets, including those for audio question answering. Like Qwen-Audio, it leverages proprietary LLMs to generate question–answer pairs grounded in the captions of existing audio corpora. LTU [4] constructs a large-scale audio question answering dataset, covering both closed-ended and open-ended formats. The closed-ended set includes tasks such as sound classification, audio captioning, and temporal analysis. Questions are generated using proprietary LLMs, while answers follow a fixed structure derived from rule-based algorithms. For open-ended data, which is more expensive and labor-intensive to create manually, LTU uses a rule-based pipeline known as audio instruction generation. This method inputs metadata such as audio captions and sound event tags into carefully designed prompts, allowing proprietary LLMs to produce diverse and contextually appropriate audio QA pairs.
On the other hand, a distinct strategy for developing speech- or audio-aware large language models focuses on leveraging the backbone LLM itself to generate synthetic data for alignment. These alignment data include general response-style texts conditioned on speech/audio, captions describing speech/audio under textual conditions, or continuation responses of speech/audio content based on textual prompts. Several previous works [16, 14, 17, 19, 18] have adopted this approach. Unlike studies that rely on externally generated task-specific question-answer pairs [1, 2, 3, 4, 5], these works utilized a backbone language model to generate general-purpose alignment data directly, enabling a more streamlined and self-contained approach to integrating speech/audio with language models. DeSTA [15] found that the approach leveraged speech captioning to bridge the gap between speech and text modalities, while DeSTA2 [14] further discovered that generating captions using the backbone LLM was more efficient. AudioChatLlama [16] and DiVA [19] used the backbone language model to generate responses conditioned on textual transcriptions. BLSP [17] further extended this idea by generating continuations based on transcriptions, while BLSP-Emo [18] incorporated additional paralinguistic information like emotion to produce more expressive outputs. DeSTA2 [14] further extended the use of paralinguistic cues, such as emotion, accent, and gender, to enhance the expressiveness of generated speech captions.
Inspired by these methods and motivated by the lack of research applying this strategy to general audio settings, our work investigates how to bootstrap audio–language alignment through synthetic data generation using a backbone LLM. We further explore various extensions of this idea to evaluate its generality and effectiveness across different scenarios. For example, we exploit the flexibility and scalability of this framework to introduce negative samples and multi-audio scenario into the training process. This article provides a practical way to effectively train audio-aware LLMs without the need to manually collect or rely on proprietary models to generate a large amount of task-specific question-answer data.
III Method
III-A Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs
During the data construction phase, we aim to generate audio-language pairs while ensuring minimal textual discrepancies between the training data and the backbone LLM used in the ALLM. This design allows the model to maintain consistency in textual representation without modifying the backbone LLM’s parameters during subsequent training.
The rationale behind generating audio-language pairs using the backbone LLM is to enable the model to acquire audio understanding capabilities solely through an audio modality adapter. This approach ensures that the audio-aware large language model aligns its comprehension of audio with the text-based metadata information understood by the backbone LLM, without requiring adjustments to the backbone LLM itself.
Figure 1: The left diagram illustrates the Data Construction Stage, where a backbone LLM generates audio-language aligned descriptions for audio clips. The right diagram represents the Training Stage, where an audio modality adapter learns to align audio inputs while the backbone LLM remains frozen. In the single-audio scenario, the process follows only the orange arrows, whereas in the multi-audio scenario, both orange and blue arrows are involved, indicating the simultaneous alignment of multiple audio inputs.
Next, we provide a detailed explanation of how LLMs are leveraged to generate audio-language alignment data. While LLMs can not directly process audio, they excel at understanding and analyzing textual metadata associated with audio, such as sound events and audio captions. To construct audio-language alignment data, we utilize the metadata from original audio datasets, which typically include two common types of annotations: human-annotated captions (${D}_{{caption}}$) and sound event tags (${D}_{{tag}}$). To further minimize textual discrepancies and eliminate the need for manually designing task-specific question-answer pairs for formatting training data, we introduce a structured format called the seed prompt (${P}_{seed}$). The seed prompt is derived from the original dataset’s annotations and can take the form of either audio captions or sound event tags, depending on the available metadata. We then append generation prompts (${P}_{gen}$) to the seed prompt and feed them into the LLM, prompting it to generate descriptions aligned with the audio content. There are three types of generation prompts:
Positive Samples Generation Prompt (${P}_{{pos}}$) aims to generate descriptions of sound events that actually occur in the audio. For example, Replay the audio.
Negative Samples Generation Prompt (${P}_{{neg}}$) aims to generate descriptions of sound events that are not present in the audio. For example, Identify sounds that are absent as contrasting examples.
Combined Samples Generation Prompt (${P}_{{comb}}$) integrates both of the above prompts. It aims to generate descriptions of both the sound events that are present and those that are absent in the audio. For example, Replay the audio and identify sounds that are absent as contrasting examples.
In summary, the final input prompt (${P}_{{final}}$) to the LLM can be expressed in Equation 1 and the overall pipeline is illustrated in Figure 1.
$P_{final}=$ $\,[\textit{Begin of audio}]\,P_{seed}\,[\textit{End of audio}]\,P_{gen}$ (1)
$P_{seed}\in\{D_{caption},D_{tag}\}$
$P_{gen}\in\{P_{pos},P_{neg},P_{comb}\}.$
In Equation 1, [Begin of audio] and [End of audio] are expressed in natural language format, with the purpose of indicating to the model that the information between them represents the content from the audio. When training audio-aware large language models, the position of ${P}_{{seed}}$ will be replaced with corresponding audio representation. These generated responses to the input context are then used as our training targets. Actual examples are demonstrated in Table I. In this work, we refer to the whole data generation process in this subsection as Bootstrapping Audio-Language Alignment via Synthetic Data Generation from Backbone LLMs (BALSa).
TABLE I: Examples produced under different generation prompts
Input Audio:
Audio Caption: A woman talks nearby as water pours.
Positive Samples Generation Prompt ($\bm{P}_{\bm{pos}}$):
Replay the audio.
Response:
Water pouring sounds, woman talking in the background.
Negative Samples Generation Prompt ($\bm{P}_{\bm{neg}}$):
Identify sounds that are absent as contrasting examples.
Response:
Based on the provided audio, here are some specific sound events that are not present in the audio:
1. A car driving by   2. Birds chirping   3. A dog barking
Combined Samples Generation Prompt ($\bm{P}_{\bm{comb}}$):
Replay the audio and identify sounds that are absent as contrasting examples.
Response:
Specific sound events detected in the provided audio:
1. A woman’s voice   2. Water pouring
Contrastive examples of specific sound events not present in the provided audio:
1. A car driving by   2. A dog barking   3. The door bang
III-B BALSa-MA: Bootstrapping Multi-Audio Learning via Difference Comparison and Joint Captioning
BALSa can be extended to multi-audio scenarios, enabling the backbone LLM to both generate comparative explanations of differences between audio samples and independently caption each sample within the same response for training purposes. The objective is to move beyond single-audio alignment by introducing a multi-audio learning framework that incorporates both discrimination-based comparison and parallel description-based captioning. In this process, the model must first independently comprehend each audio sample before either distinguishing their differences or simultaneously generating separate captions for both audio samples within a unified response. This dual-task formulation increases the learning complexity compared to single-audio alignment. We refer to this multi-audio extended approach as BALSa-MA (Multi-Audio).
To implement this, we extend the generation prompt to accommodate multiple audio inputs. Here, we consider a two-audio scenario, denoted as first audio and second audio. The seed prompt also can be either audio captions (${D}_{{caption}}$) or sound event tags (${D}_{{tag}}$), depending on the annotation format of the original dataset. We denote them as ${P}_{{seed,1}}$ and ${P}_{{seed,2}}$, respectively. Next, we append the generation prompts (${P}_{gen}$) to both seed prompts and feed them into the LLM, enabling it to generate discrimination-based descriptions that compare the two audio samples.
In this study, we employ the following generation prompt: Discrimination Generation Prompt (${P}_{{diff}}$), which generates comparative descriptions highlighting the differences between two audio samples. For instance, Explain the difference between the audio clips.
The Both Audio Captioning Prompt (${P}_{{cap,both}}$), which generates separate captions for each audio sample, providing distinct descriptions for both inputs within a single response.
In summary, the final input prompt (${P}_{{final}}$) to the LLM can be expressed in Equation 2 and the overall pipeline is illustrated in Figure 1.
[…]
$P_{final}$ $=P_{final,1}\,||\,P_{final,2}\,||\,P_{gen}$
$P_{seed}$$\in\{D_{caption},D_{tag}\}$
$P_{gen}$$\in\{P_{diff},P_{cap,both}\}.$
In Equation 2, $\parallel$ denotes concatenation, while [Begin of audio] and [End of audio] are expressed in natural language format to serve two main purposes. First, they indicate to the model that the enclosed content corresponds to the audio. Second, they provide flexibility for adaptation to multiple-audio formats. When training audio-aware large language models, the positions of ${P}_{{seed,1}}$ and ${P}_{{seed,2}}$ will be replaced with the corresponding audio representations. The generated responses based on the input context are then used as training targets. Actual examples are presented in Table II.
TABLE II: Examples corresponding to different generation prompts.
Input Audio:
Audio1 Caption: A woman talks nearby while water pours.
Audio2 Caption: A dog barks and a man speaks.
Discrimination Generation Prompt ($\bm{P}_{\bm{pos}}$):
Explain the difference between two audio.
Response:
Audio1 features a woman speaking with the background sound of pouring water, while Audio2 contains a dog barking alongside a man’s speech, differing in both speaker identity and background sound events.
Both Audio Captioning Prompt ($\bm{P}_{\bm{comb}}$):
Repeat the both audio.
Response:
Audio1: The sound of water pouring accompanies a woman speaking nearby.
Audio2: A man talks while a dog barks in the background.
III-C Model Design and Training Approach
Figure 1 illustrates the overall framework. The model comprises three primary components: an audio encoder, a backbone large language model, and an audio modality adapter.
Audio Encoder. We employ Whisper [37], a foundation model, as the audio encoder. Previous studies [38] have demonstrated Whisper’s strong performance across various audio-related tasks, despite its original design being primarily for automatic speech recognition and speech translation. Since Whisper follows an encoder-decoder architecture, we utilize only its encoder component. To preserve the advantages of the pre-trained model, the parameters of the audio encoder remain frozen.
Backbone Large Language Model. This study adopts LLaMA-3.1-8B-Instruct [30], an instruction-tuned model, as the backbone large language model. To maintain its original text-processing capabilities, we do not apply LoRA or fine-tune any of its parameters.
Audio Modality Adapter. The only trainable component is the audio modality adapter, which is randomly initialized. This adapter projects the output representations extracted by the audio encoder into the input space of the backbone large language model. Specifically, we employ a Qformer [39] architecture to extract audio features from the intermediate layers of the Whisper encoder. These extracted features are aggregated using a weighted sum with learnable weights, followed by a linear projection layer that aligns the aggregated features with the input dimensions of the backbone large language model. We optimize the entire architecture end-to-end with a next-token prediction loss, using the model’s own generated responses as learning targets. To improve efficiency, we freeze the parameters of both the backbone LLM and the audio encoder. Instead, we adopt a lightweight adapter-based training strategy [40, 15, 17, 18, 19], in which only the modality adapter is trained. This design not only conserves computational resources but also preserves the LLM’s original linguistic knowledge while enabling it to process audio through the adapter.
IV Experimental Setup
IV-A Training Datasets
The training datasets include AudioSet-20K [41], AudioCaps [42], FSD50K [43], MACS [44], ESC-50 [45], UrbanSound8K [46], Clotho [47], and VocalSound [48]. Among these datasets, AudioCaps, Clotho, and MACS provide ground-truth labels in the form of human-annotated audio captions, while the rest use sound event tags. Table III summarizes the total duration of the training datasets used in our work and compares it with the training data durations reported in previous studies [1, 2, 4, 5, 3]. The “Duration” column represents the original total duration of unique audio in each dataset, measured in hours and excluding any duplicate segments. The “Equi. Duration” column accounts for the effect of generating multiple aligned text descriptions per audio clip, leading to an increased effective training duration, also measured in hours. The blue numbers indicate the amount of data used in the multi-audio scenario. The number of samples represents the total number of data instances used during training.
TABLE III: Summary of training datasets.
Dataset Duration Equi. Duration # Samples
AudioSet-20K[41] 60 121 42K
AudioCaps[42] 138 276+269 96K+49K
Clotho[47] 18 36+160 27K+13K
MACS[44] 11 22 33K
FSD50K[43] 56 113+75 56K+19K
ESC50[45] 3 6+28 4K+10K
UrbanSound8K[46] 10 19 17K
VocalSound[48] 20 40 34K
Total (Ours) 316 1164 400K
Qwen2-Audio[1, 2] - 10000 -
SALMONN[3] - 1044 370K
LTU[4], LTU-AS[5] - 1784 5682K
IV-B Model Selection and Training Setups
Our framework integrates two key foundation models: Llama-3.1-8B-Instruct333huggingface.co/meta-llama/Llama-3.1-8B-Instruct[30] and the compact Whisper-small444huggingface.co/openai/whisper-small[37], which contains about 240 million parameters. To connect these components, followed by [40, 15, 17, 18, 19], we developed a specialized modality adapter, featuring a two-layer transformer decoder architecture. This adapter uses 64 specialized vectors to capture and process audio features from the encoder’s hidden states. In our implementation strategy, we maintain the original architecture of both LLaMA and Whisper unchanged. Instead, we focused our training efforts solely on the modality adapter, which introduces approximately 22 million trainable parameters. The system processes information by first converting text instructions through LLaMA’s embedding layer, then combining them with the processed audio features.
To enhance the model’s ability to handle multi-audio inputs, a progressive learning strategy is adopted, where the model is first trained on single-audio tasks before transitioning to multi-audio scenarios, following a curriculum learning approach. For the technical implementation, we built our system using the Hugging Face Transformers[49] framework. Following our progressive learning strategy, the training process consisted of two stages: first, we pretrained the model on single-audio tasks for five epochs to establish a strong baseline; second, we fine-tuned it under both positive and negative data settings on multi-audio data for two epochs to enhance its ability to process multiple inputs. We used the Adam optimization algorithm with a cosine annealing learning schedule and 2,000 warm-up steps. All experiments were conducted on two NVIDIA RTX 6000 Ada Generation GPUs, with a global batch size of 16 and a learning rate of 1e-4.
IV-C Baseline Models and Evaluation Setups
In this study, we incorporate widely used, well-documented, and fully open-source models as baselines for comparison. These include Qwen-Audio-Chat [1], Qwen2-Audio-Instruct [2], SALMONN [3] (including 7B and 13B versions), LTU [4], and LTU-AS [5], as well as proprietary commercial models like Gemini-Pro-1.5 [36]. Since most ALLMs are instruction-tuned for open-ended responses [1, 2, 3, 4, 5], they generate answers based on user queries rather than selecting from predefined choices. To align with this design, we extract relevant answers from model outputs using a structured evaluation approach. All experiments adopt a consistent greedy decoding strategy with a maximum output length of 512.
To facilitate the extraction of relevant answers during evaluation, we implement robust regular expressions to extract key information from model responses. The extracted content is then matched to the provided options via string-based comparison. To minimize potential biases, the answer options in the original question are randomized before evaluation. To ensure concise questions, we randomly select three incorrect options along with the correct answer and then shuffle their order in the multiple-choice format.
Once the relevant answers are extracted and mapped to the given options, we proceed with the evaluation process using standardized metrics. If a model’s response cannot be parsed using the aforementioned approach, it is considered incorrect. Our evaluation strategy is structured as follows. We report the weighted F1 score for multi-class classification tasks. For other tasks with their own evaluation protocols, we follow established methodologies [13, 12, 50, 51, 52, 53]. In our experiment results, bold text indicates the best performance, while underlined text represents the second-best. We further detail the evaluation methods for each benchmark.
Audio Question Answering Benchmark. In ClothoAQA, which includes both binary and non-binary classification tasks, we report accuracy as the primary metric.
Audio Reasoning Benchmark. We report weighted F1 score for multi-class classification tasks. For Synonym-Hypernym Test, we also report weighted precision and recall. In addition, the original MMAU benchmark uses micro-averaged accuracy as its evaluation metric. The SAKURA test set follows a multiple-choice format, and we evaluate model performance using three key metrics. We report overall accuracy across all questions, as well as separate accuracy scores for single-hop and multi-hop questions.
Audio Hallucination Benchmark. Following previous studies [13], we compute overall accuracy and F1 scores for questions where the correct answer is “Yes” or “No”. As an additional reference, we also report the proportion of cases where the model responds with “Yes”.
Instruction-Following Benchmark. We report the accuracy of each instruction-following task, the overall average accuracy, and the performance gap between the evaluated model and its backbone LLM.
IV-D Ablation Study
First, in the single-audio training scenario, to ensure a fair evaluation of contrastive-like methods using BALSa, we conducted the following experiments to investigate the role of negative samples.
Positive-only Training Data: The training set contains only positive samples, with a total of $2N$ data points.
Positive and Negative Training Data: The training set includes both positive and negative samples, each with $N$ data points, resulting in a total of $2N$ data points.
Combined Training Data: The training set consists of combined samples, as described in Section III. A combined sample includes both sound events that are present and those that are absent within a single sample. In contrast, positive samples only contain present sound events, while negative samples only contain absent sound events. Under the combined setup, there are $N$ data points. Since each data point carries information about both present and absent sound events, it is equivalent to having a total of $2N$ data points.
Second, to validate that data generated by the backbone LLM outperforms data generated by an external LLM under this setting, we also examine a scenario where a stronger LLM is used to generate synthetic data via our proposed pipeline, BALSa. Specifically, we select Gemini-1.5-Pro 555gemini-1.5-pro-latest to synthesize the positive-only, positive and negative, and combined training data. The same methodology is applied, but a more advanced model is used to generate the data.
Third, we compare two approaches for generating training data for audio-language alignment: our method, which allows the backbone LLM to freely generate captions based on meta-information, and another that provides it with collected questions and asks it to generate corresponding answers. To investigate this, we leverage the questions (${Q}_{{OpenAQA}}$) from the OpenAQA dataset [5] and apply the BALSa pipeline to generate corresponding answers using the backbone LLM. Specifically, OpenAQA contains 5 million instances, from which we select a subset that aligns with the audio data used in this work. As a result, the final training dataset also consists of $2N$ data points. In summary, the final input prompt (${P}_{{final}}$) to the backbone LLM is formulated in Equation 3. This setting refers to BALSa-OpenAQA.
$P_{final}=$ $\,[\textit{Begin of audio}]\,P_{seed}\,[\textit{End of audio}]\,P_{gen}$ (3)
$P_{seed}\in\{D_{caption},D_{tag}\}$
$P_{gen}\in\{Q_{OpenAQA}\}.$
For comparison, we also directly use question-answer pairs derived from the OpenAQA dataset as our training data.
Fourth, we explore an alternative approach where, instead of using BALSa to generate captioning data, we employ pre-designed description-based prompts like ”Describe the audio.”, which is the rule-based templates. In this setting, the training targets are the corresponding ground truth captions or sound event tags of the audio. The total training dataset also consists of $2N$ data points.
Beyond the single-audio scenario, we conduct ablation experiments on the multi-audio scenario, BALSa-MA, an extended version of BALSa. In BALSa-MA, we aim to enable models to learn from multiple audio sources, including explaining the differences between two audio samples and captioning both simultaneously. For conciseness, in our experiments, we adapt BALSa-MA to a two-audio-sample setting.
To ensure a fair evaluation of different methods for audio interpretation, we design the following experiments.
Discrimination-only Training Data: The training set contains only discrimination samples, with a total of $2M$ data points.
Captioning-only Training Data: The training set contains only captioning samples, with a total of $2M$ data points.
Joint Discrimination-Captioning Training Data: This training set is a balanced combination of discrimination and captioning samples, with each type comprising 50% of the dataset, totaling $2M$ data points.
V Evaluation Benchmarks
In this work, we categorize evaluation benchmarks into three types: audio question answering benchmark, audio reasoning benchmark, and reliability and safety benchmark.
V-A Audio Question Answering Benchmark
This benchmark focuses on assessing a model’s fundamental understanding of audio. It evaluates capabilities such as sound event recognition, acoustic scene classification, temporal comprehension, and audio attribute identification. By designing diverse types of questions, this benchmark examines whether a model can accurately interpret various aspects of an audio signal and provide reasonable and correct answers based on the given auditory information. This allows us to gauge the model’s core competence in audio understanding. For this benchmark, we have collected and organized the following evaluation datasets.
Clotho-AQA: Clotho-AQA [47] is a widely used dataset designed for audio question answering (AQA) tasks, where models generate natural language answers based on audio signals and corresponding questions. The dataset is constructed through crowd-sourcing, with questions collected for each audio file and answers provided by multiple annotators. To ensure data reliability, only questions where all annotators unanimously agree on the answer are included. The final dataset consists of 789 instances.
Nonspeech7k AQA: Nonspeech7k [54] is a dataset of human non-speech sounds collected from real-life environments, covering seven sound categories: breath, cough, cry, laugh, scream, sneeze, and yawn. We utilize all 725 audio instances from the test split of the Nonspeech7k dataset for evaluation. To adapt the dataset to a question-answering (QA) format, we manually designed general questions such as: Classify the given sound as exactly one of: breath, cough, cry, laugh, scream, sneeze, or yawn. This ensures that the model follows explicit instructions when answering the questions.
CochlScene AQA: CochlScene [55] is a crowd-sourced dataset designed for acoustic scene classification, covering 13 different acoustic scenes. The primary goal of this task is to determine the environment in which a given audio recording was captured based on its acoustic characteristics. For evaluation, we sample 100 audio clips from each category in the test split, resulting in a total of 1,300 test instances. We converted the dataset into a question-answering (QA) format through the manual design of general questions like Which acoustic scene does this audio clip belong to? The answer must be one of the four given options: Residential Area, Restaurant, Subway Station, Car.
EDANSA-2019 AQA: The Ecoacoustic Dataset from Arctic North Slope Alaska (EDANSA-2019) [56] is an ecoacoustic dataset collected from the North Slope of Alaska during the summer of 2019. The dataset contains over 27 hours of labeled audio. Since this dataset was recorded in the Arctic, most existing models may not have been trained on similar data. Publicly available training datasets rarely include such recordings, making EDANSA-2019 an out-of-distribution (OOD) dataset. This poses a challenge for models, as they may struggle to generalize to these unfamiliar acoustic environments.
For evaluation, we selected sound events with a sufficient number of samples, covering fly, bird, woof, wind, rain, aircraft, and machine. Due to data limitations, fly (62 instances) and woof (15 instances) have fewer samples, while all other categories contain 200 randomly selected audio clips. In total, the test set consists of 1,077 instances. We reformulated the dataset into a question-answering (QA) format by manually crafting general questions, for example: Select the most appropriate category for the given sound from the following options: Wind, Machine, Rain, Fly.
V-B Audio Reasoning Benchmark
This benchmark focuses on evaluating a model’s advanced reasoning capabilities in audio understanding. Beyond recognizing and interpreting audio content, models must demonstrate the ability to reason based on their understanding to make logical inferences. For this benchmark, we utilize the following dataset for this benchmark:
Synonym and Hypernym Test: A previous study [57] introduced a synonym and hypernym test to assess whether models can understand audio and its semantic connections to text. Synonyms refer to words with similar meanings, while hypernyms denote broader categories that encompass more specific terms. These relationships enable the construction of text-based prompts that test a model’s understanding of similarity and hierarchy in audio reasoning. For example, given an audio clip of a songbird chirping, a possible question could be: Is the sound from an object that is a type of chordate? The expected answer is yes.
This benchmark evaluates whether audio-aware large language models genuinely comprehend audio content and semantic relationships, or if they merely rely on statistical associations between audio, keywords, and captions. We reproduce the evaluation setup and data selection from the original study, sourcing data from the AudioCaps dataset, resulting in 636 evaluation instances. To further analyze the model’s behavior, we separately compute weighted precision, recall, and F1 scores for questions where the correct answer is yes and no.
Audio Entailment: A previous study [58] introduced the task of audio entailment to assess a model’s ability to perform deductive reasoning. This task evaluates whether a textual description (hypothesis) can be logically inferred from an audio recording (premise). The conclusion falls into one of three categories: entailment, neutral, or contradiction, depending on the available evidence. Entailment occurs when the audio recording provides sufficient evidence to affirm the truth of the hypothesis. Neutral applies when the audio recording does not contain enough information to either confirm or deny the hypothesis. Contradiction is determined when the audio recording offers clear evidence that the hypothesis is false.
We format this task as a multiple-choice question, requiring the model to identify the correct conclusion type based on the given prompt. For example, given an audio clip, the model is asked to select the hypothesis that corresponds to entailment: Based on what you hear, which of the following hypotheses is entailment? (A) A quiet room with no movement or wind. (B) A motorcycle is overtaking cars on a windy road. (C) Vehicles are moving on a road with wind noise present. In this setup, the order of the answer choices is randomly shuffled for each instance. The dataset sources audio recordings from Clotho [47] and AudioCaps [42]. For evaluation, we report the overall accuracy.
SAKURA: This benchmark [50] evaluates multi-hop reasoning in audio understanding. We use the SAKURA test set from the audio track for this evaluation. In this benchmark, each audio clip is paired with two types of questions: single-hop and multi-hop. Single-hop questions directly ask about attributes of the sound event in the audio, such as Out of the animals listed below, can you identify the one that most closely matches the sound in the audio? In contrast, multi-hop questions require multiple steps of reasoning, combining fundamental audio understanding with text-based inference to arrive at the correct answer. For example, a multi-hop question might be: Taking into account the sound from the recording and the natural tendencies of this type of animal making the sound, which of the following behaviors is most representative of how it typically acts? (a) Wagging its tail (b) Jumping between branches (c) Hiding in shells (d) Flapping its wings.
MMAU: This is a massive multi-task audio understanding and reasoning benchmark [51] designed to evaluate multimodal audio understanding models on tasks requiring expert-level knowledge and complex reasoning. In this study, we focus on assessing model performance specifically along the audio dimension, using the test-mini split for evaluation. The evaluation settings follow the original benchmark’s guidelines.
V-C Reliability and Safety Benchmark
The Reliable and Safety Benchmark focuses on evaluating the trustworthiness and safety of audio-aware large language models. We select the following evaluation benchmark.
Audio Hallucination: Our previous work [13] introduced benchmarks to evaluate object hallucination in audio—a phenomenon where models incorrectly detect or infer non-existent sound events. In this task, given an audio clip, the model is tested on whether a specific sound event is present. Inspired by Polling-based Object Probing Evaluation (POPE) [59], our prior work adopted a similar methodology to evaluate ALLMs. We formulate object hallucination detection as a binary classification task, prompting ALLMs to respond with “Yes” or “No”. To assess model robustness, we design four different prompts that ask whether a particular sound is present, including variations such as Is there a sound of [object]?, Does the audio contain the sound of [object]?, Have you noticed the sound of [object]?, and Can you hear the sound of [object]?
This benchmark is constructed using AudioCaps, sampling audio clips that contain more than three sound events. It consists of 407 distinct audio clips, where we treat all ground truth labels as positive samples. To introduce negative samples, we apply a random negative sampling strategy, selecting three objects that are not present in the given audio. In addition, in order to further analyze the model’s behavior, we separately compute precision, recall, and F1 scores for questions where the correct answer is yes and no.
Instruction Following Ability Evaluation: Previous work [12] has assessed instruction-following capabilities and quantified catastrophic forgetting in audio-aware large language models. The findings highlight that most audio-aware large language models struggle with even basic instruction following, exhibiting significant performance degradation compared to text-based large language models. To evaluate the model’s instruction-following ability, we utilize both close-ended and open-ended questions. Close-ended questions require the model to adhere to specific formatting constraints, ensuring strict compliance with instructions. In contrast, open-ended questions allow for more flexibility, evaluating the model’s ability to follow complex and creative instructions. We report the instruction following rate (IFrate) and forgetting rate ($\Delta$), which is the relative IFrate difference between ALLMs and their LLM counterpart. A smaller forgetting rate indicates a lower degree of forgetting in ALLMs. Note that since we cannot directly access proprietary models like Gemini, we are unable to calculate the forgetting rate. Similarly, for the Qwen-Audio models [1, 2], which are finetuned from the pretrained Qwen-Audio base model [1], it is not feasible to compute the IFrate difference between ALLMs and their text-only LLM.
VI Experimental Results
VI-A Performance of BALSa
TABLE IV: Evaluation results for the audio question answering and audio hallucination benchmark. F1 (Y) and F1 (N) represent the F1 scores for cases where the correct answer is yes and no, respectively. F1 (W) denotes the weighted F1 score. Acc refers to accuracy, and Yes indicates the percentage of responses where the model answers yes. (Unit: %)
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination
Models F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes
Qwen-Audio-Chat[1] 34.50 74.90 55.58 34.73 75.25 52.97 64.11 81.04
Qwen2-Audio-Instruct[2] 41.40 75.55 60.16 54.84 72.89 54.44 63.67 75.39
SALMONN-7B[3] 26.36 76.56 37.99 78.29 70.06 26.48 48.27 92.14
SALMONN-13B[3] 29.96 83.99 31.06 77.24 76.94 60.07 68.50 76.78
LTU[4] 38.71 67.10 14.91 5.79 78.27 73.65 75.96 59.97
LTU-AS[5] 23.16 56.77 20.30 8.06 51.68 45.49 48.59 56.02
Gemini-1.5-Pro[36] 17.46 68.43 36.18 47.01 65.86 70.95 68.41 42.02
Ours (BALSa, only-positive) 45.57 75.11 39.28 54.17 73.05 55.15 64.10 74.81
Ours (BALSa, positive-negative) 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69
Ours (BALSa, combined) 48.56 80.93 42.48 52.17 77.21 71.23 74.22 61.58
Ours (BALSa-MA, discrimination) 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29
Ours (BALSa-MA, captioning) 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01
In Table IV, compared to other publicly available baseline models, our proposed audio-aware large language models trained using the BALSa method (combined, positive-only, and positive-negative) achieve performance comparable to other public baseline models in audio question answering benchmarks. For instance, in ClothoAQA, our models perform on par with Qwen2-Audio-Instruct and SALMONN-7B. In NonSpeech AQA, our models attain results comparable to Qwen2-Audio-Instruct and Qwen-Audio-Chat. In CochlScene AQA, our models outperform SALMONN-7B and SALMONN-13B. Compared to Gemini-1.5-Pro, a proprietary model, our method remains competitive, leading by 6% to 25% on these evaluation benchmarks. While SALMONN-13B achieves the best performance on ClothoAQA and NonSpeech AQA, our models surpass it in other evaluation benchmarks. Despite falling behind Qwen-Audio-Chat and Qwen2-Audio-Instruct in tasks such as CochlScene AQA, our models demonstrate well-rounded performance across various benchmarks. These results highlight the effectiveness of the BALSa framework. By leveraging synthetic data to enhance audio-language alignment, BALSa delivers competitive results on traditional audio question answering benchmarks, often surpassing other baselines while maintaining consistent performance across different benchmarks. Notably, our models achieve these results using only 12% of the equivalent training duration required by the widely adopted and powerful Qwen2-Audio-Instruct model, as shown in Table III.
The EDANSA AQA tasks serve as zero-shot benchmarks, evaluating models on out-of-domain audio that is not encountered during training. Our models, along with SALMONN and LTU-AS, are assessed under these conditions. Compared to other baselines, our method demonstrates a significant performance advantage. The results from this benchmark suggest that our approach exhibits a certain level of adaptability when applied to out-of-distribution data. Note that the Qwen series does not disclose its training datasets, thereby making it impossible to determine whether its results are obtained under a true zero-shot setting.
In audio hallucination tasks, incorporating negative samples through the BALSa framework proves to be beneficial. After adding negative samples, the positive-negative setting improves the weighted F1 score by 13%, representing an approximate 20% increase compared to the positive-only setting. This suggests that a contrastive-like approach effectively help mitigate audio hallucinations while preserving performance on audio question answering. Notably, in audio hallucination tasks, many baseline models struggle to distinguish nonexistent sounds when the ground truth is no, frequently misidentifying them as present. In fact, most baseline models tend to answer yes regardless of whether the sound exists. This issue raises concerns about model reliability in real-world applications. Even models that perform well on other evaluation benchmarks, such as Qwen2-Audio-Instruct and Qwen-Audio-Chat, exhibit weaker performance in hallucination detection. Our proposed method leverages BALSa to generate negative samples. The addition of these samples, while simple, directly and effectively enhances performance on the audio hallucination benchmark. At the same time, it maintains performance on other evaluation benchmarks and, in some cases, even leads to slight improvements. This demonstrates that our approach not only mitigates hallucinations but also achieves a strong level of performance in fundamental audio understanding. Regarding the incorporation of negative samples, our experimental results indicate that separating positive and negative samples yields better performance compared to combining them into a single sample.
TABLE V: Evaluation results for the audio reasoning and instruction-following benchmark. P represents the weighted precision, R denotes the weighted recall, and F1 refers to the weighted F1 score. Acc indicates overall accuracy, while S-Accuracy and M-Accuracy measure accuracy for single-hop and multi-hop questions, respectively. Acc (C) and Acc (A) represent accuracy when the source datasets are Clotho and AudioCaps, respectively. (Unit: %)
Synonym-Hypernym SAKURA MMAU Audio Entailment Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc (C) Acc (A) Close Open IFRate $\Delta$
Qwen-Audio-Chat[1] 86.26 79.88 79.93 78.00 89.20 67.26 38.74 36.65 34.04 10.93 56.00 32.00 -
Qwen2-Audio-Instruct[2] 82.52 68.71 67.67 76.90 92.00 62.83 57.96 42.97 46.19 41.59 32.00 47.11 -
SALMONN-7B[3] 61.72 54.71 54.04 52.40 62.40 44.23 51.95 33.24 34.10 63.02 46.00 54.51 -16.27
SALMONN-13B[3] 76.55 61.16 59.04 58.30 70.00 46.86 48.95 35.25 30.81 37.41 61.25 36.89 -34.53
LTU[4] 73.82 72.64 70.47 51.10 62.80 38.85 22.22 18.88 14.42 9.97 38.75 24.36 -62.58
LTU-AS[5] 47.91 50.63 48.71 24.70 30.40 19.74 16.82 26.95 28.64 28.83 47.75 29.19 -41.18
Gemini-1.5-Pro[36] 73.84 73.42 73.35 44.70 41.20 55.34 48.65 23.32 22.87 99.14 89.50 94.32 -
Ours (BALSa, only-positive) 82.15 78.77 78.98 71.60 79.20 64.14 55.26 50.88 52.41 90.57 90.25 90.41 -0.68
Ours (BALSa, positive-negative) 77.94 77.04 77.25 70.70 76.40 65.71 55.26 56.30 57.16 90.78 91.25 91.02 -0.02
Ours (BALSa, combined) 75.12 75.31 74.76 65.00 68.40 61.99 54.95 46.92 49.00 88.32 91.50 89.91 -1.23
Ours (BALSa-MA, discrimination) 81.76 80.50 80.55 79.80 86.00 73.49 56.76 54.67 55.66 88.85 87.25 88.05 -3.27
Ours (BALSa-MA, captioning) 82.99 82.86 82.89 80.30 88.40 71.72 57.06 56.40 56.59 89.07 84.25 86.66 -4.80
In the audio reasoning benchmark, which is demonstrated in Table V, models are required to rely on their fundamental audio understanding and apply reasoning skills to determine the correct answer. One of the key objectives of audio-language alignment is to integrate audio information, which belongs to a different modality than text, into the text space of large language models. However, few studies [57] have systematically and explicitly explored this issue. As an initial exploration, we build upon previous work [57, 50] and evaluate our model, along with other baselines, on several public benchmarks, including Synonym-Hypernym Test and SAKURA [50]. These evaluation aim to determine whether these models can genuinely comprehend audio content and utilize it for reasoning about semantic relationships. In the Synonym-Hypernym Test, our proposed models achieve impressive performance, demonstrating stronger and more robust audio-text alignment compared to other baselines. The only-positive setting attains an F1 score of 78.98, trailing Qwen-Audio-Chat by only a marginal difference of 1.0 point, while surpassing all other public baseline models by 8% to 31%. Although Qwen-Audio-Chat achieves comparable performance, our method requires only 12% of their dataset’s total hours and does not involve any modifications to the audio foundation model or the backbone large language model.
Furthermore, SAKURA [50] is a multi-hop reasoning benchmark designed to assess whether a model can first comprehend information within audio and then perform multi-hop reasoning based on that information. If a model merely memorizes the correspondence between text and audio during training, it may fail to flexibly apply its learned knowledge in reasoning tasks. Although our models initially achieve lower overall accuracy compared to Qwen2-Audio-Instruct, they exhibit comparable performance in multi-hop question accuracy. Our analysis suggests that the weaker performance on the SAKURA benchmark is primarily due to lower accuracy in single-hop questions, which in turn affects performance on multi-hop questions. To address this limitation, we later incorporate multi-audio training, effectively enhancing the model’s performance.
In audio entailment tasks, models are evaluated on their ability to perform deductive reasoning for audio understanding. Experimental results show that our proposed models, including both the only-positive and positive and negative settings, outperform other baseline models by at least 10% in audio entailment (Clotho). While the performance gap in audio entailment (AudioCaps) is smaller, our models still achieve better results compared to other baselines. In MMAU, compared to the state-of-the-art model on the sound track of MMAU test-mini, Qwen2-Audio-Instruct, our models lag by only 2.7%, demonstrating a comparable level of performance. Furthermore, across all other comparisons, our models surpass other baselines by at least 3% to 20%. The results of audio reasoning benchmark indicate that BALSa achieves an audio-language alignment performance comparable to other models. Notably, our approach does not require constructing an extensive task-specific question-answering dataset, making it a practical and feasible alternative for future training of audio-language alignment capabilities.
For the evaluation of instruction-following abilities, excluding the proprietary model Gemini-1.5-Pro, our models achieve the highest instruction-following rate and the lowest forgetting rate. Note that the performance of the baseline models is sourced from [12]. This benchmark reassesses the fundamental aspects of developing ALLMs by separately examining their instruction-following capabilities and their understanding of audio content. Models with strong audio understanding, such as Qwen2-Audio-Instruct and SALMONN-13B, struggle with even basic instruction-following. They fail to retain their textual capabilities after audio-language alignment training. This issue is evident in Qwen2-Audio-Instruct, which achieves an instruction-following rate of only 47.1%, while SALMONN-13B suffers from a forgetting rate as high as 34%. During training, if chasing audio-language alignment leads to a significant loss of existing abilities, such as following basic instructions, it might reduce the model’s effectiveness in real-world applications.
In summary, our experimental results demonstrate that our models achieve a balance between audio understanding, reasoning, and fundamental instruction-following abilities. They effectively combine strong audio understanding capabilities with robust instruction-following performance.
VI-B Comparison of Different Data Generation Approaches
TABLE VI: Evaluation results of our ablation study on the audio question answering and audio hallucination benchmark. (Unit: %)
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination
Models F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes
Ours (Gemini, only-positive) 38.39 46.58 31.21 40.94 66.58 3.35 34.96 98.61
Ours (Gemini, positive-negative) 16.09 52.98 29.26 18.99 63.92 32.90 48.41 80.06
Ours (Gemini, combined) 48.53 66.81 35.68 26.75 58.76 66.89 62.83 39.07
Ours (Rule-based Template) 22.40 38.57 15.19 22.29 66.14 1.24 33.69 77.44
Ours (OpenAQA) 18.76 68.70 33.42 46.39 78.89 71.90 75.40 63.88
Ours (BALSa-OpenAQA) 45.77 78.31 41.69 56.22 78.67 77.65 78.16 52.33
Ours (BALSa, positive-negative) 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69
Ours (BALSa-MA, discrimination) 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29
Ours (BALSa-MA, captioning) 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01
Ours (BALSa-MA, joint) 48.02 74.82 46.03 70.73 84.19 83.78 83.98 51.27
TABLE VII: Evaluation results of our ablation study on the audio reasoning and instruction-following benchmark. (Unit: %)
Synonym-Hypernym SAKURA MMAU Audio Entailment Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc (C) Acc (A) Close Open IFRate $\Delta$
Ours (Gemini, only-positive) 31.94 36.32 32.88 47.10 54.60 40.66 45.65 35.82 33.91 52.73 63.00 57.87 -
Ours (Gemini, positive-negative) 48.67 48.59 48.23 43.00 48.00 38.75 39.64 35.02 23.92 77.92 77.50 77.71 -
Ours (Gemini, combined) 48.72 47.64 47.93 49.90 49.80 51.00 51.65 32.44 32.25 88.85 75.75 82.30 -
Ours (Rule-based Template) 16.26 23.11 18.78 31.30 54.80 7.30 10.51 3.70 0.73 0.00 29.25 14.63 -83.93
Ours (OpenAQA) 66.21 66.98 65.99 33.70 34.00 39.41 51.05 45.74 44.11 57.88 65.75 61.81 -32.10
Ours (BALSa-OpenAQA) 74.42 73.74 73.92 64.10 67.40 66.47 53.75 54.23 56.14 88.75 90.00 89.37 -1.82
Ours (BALSa, positive-negative) 77.94 77.04 77.25 70.70 76.40 65.71 55.26 56.30 57.16 90.78 91.25 91.02 -0.02
Ours (BALSa-MA, discrimination) 81.76 80.50 80.55 79.80 86.00 73.49 56.76 54.67 55.66 88.85 87.25 88.05 -3.27
Ours (BALSa-MA, captioning) 82.99 82.86 82.89 80.30 88.40 71.72 57.06 56.40 56.59 89.07 84.25 86.66 -4.80
Ours (BALSa-MA, joint) 78.43 78.46 78.28 79.30 86.00 72.79 56.16 48.92 54.80 90.25 88.25 89.25 -1.96
While our models demonstrate a balanced integration of audio understanding, reasoning, and instruction-following abilities, it is essential to analyze the factors that contribute to this performance. To further investigate the role of different components, we conduct an ablation study to examine their impact on model effectiveness. First, we replace the original backbone LLM with a stronger LLM like Gemini [36], while applying the same BALSa pipeline to generate training data. As shown in Table VI and Table VII, we observe that across all evaluation benchmarks, performance significantly declines regardless of whether the setting is positive-only, positive-negative, or combined. This result highlights the importance of using the original backbone LLM for generating audio-language alignment training data.
Second, we explore an alternative approach by providing the model with a set of pre-collected questions source from OpenAQA [4] and generating corresponding answers using the BALSa pipeline. We refer to this variant as BALSa-OpenAQA. Experimental results indicate that BALSa-OpenAQA, in most cases, perform slightly worse than those trained with BALSa-generated positive and negative samples. However, the performance gap is not substantial overall. On the other hand, if we directly train the model using the original question-answer pairs provided in OpenAQA [4], rather than generating answers through the BALSa approach, the experimental results show that models trained with BALSa-OpenAQA consistently outperform those trained directly on the original OpenAQA dataset across all evaluation benchmarks. Third, we explore an approach that utilizes pre-defined rule-based templates to construct paired training data. However, models trained with this method not only perform poorly across evaluation benchmarks but also exhibit a parroting effect, where responses frequently fail to follow instructions and instead generate only audio captions.
Based on the experimental results, we report the following findings and observations: The BALSa framework can generate audio-language alignment training data using only simple generation prompts, such as basic captioning-based instructions like Repeat the audio. If additional pre-collected question-answer data like OpenAQA [4] is available, it can also be processed through the BALSa pipeline to generate training data, with the pre-collected questions serving to increase the diversity of generation prompts. However, this approach requires additional data collection and preparation efforts. The finding in this article further reinforces the established practice within the community of using the original backbone LLM for generating training data [14], underscoring its essential role in improving model performance.
VI-C Performance of BALSa-MA
Furthermore, BALSa can be extended to multi-audio scenarios, leveraging explanations between two audio inputs to enhance the model’s audio understanding capabilities. Building on this, we extend BALSa to the multi-audio setting (BALSa-MA). Experimental results indicate that this extension leads to performance improvements across multiple evaluation benchmarks. By comparing the positive-negative setting in the single-audio scenario with the discrimination setting in the multi-audio scenario, we observe an 8% improvement in audio hallucination performance, representing an approximately 10% increase. On the fundamental audio question-answering benchmarks, our method achieves performance improvements of 8% and 15% on the EDANSA and NonSpeech tasks, respectively. On SAKURA, we observe a nearly 10% increase in overall accuracy, along with an 8% improvement specifically on multi-hop questions. In contrast, improvements in Audio Entailment are relatively smaller. Overall, we find that BALSa-MA, which enables the model to learn from two audio inputs simultaneously, enhances its basic audio understanding abilities, thereby benefiting more advanced reasoning tasks.
Following these findings, we investigate whether learning through discrimination-based training—where the model explains the differences between two audio inputs—or learning through complete description—where the model captions both audio inputs—is sufficient for improving performance. As presented in Table VI and Table VII, we observe that the performance difference between Both Captions Only (BALSa-MA, captioning) and Comparison Only (BALSa-MA, discrimination) is minimal. These findings suggest that simply learning to process two audio inputs simultaneously is sufficient to enhance the model’s overall understanding, without requiring explicit comparative training. Furthermore, we experimented with a mixed dataset, combining Both Captions Only and Discrimination-Only data in equal proportions. However, the results indicate that this hybrid approach does not provide any significant benefits. Therefore, within the BALSa-MA framework, enabling the model to jointly learn from two audio inputs is sufficient to improve its audio understanding and reasoning capabilities.
VII Conclusion
In this work, we introduce BALSa, a framework that synthesizes general-purpose caption-style alignment data for training ALLM. This framework reduces the reliance on large collections of task-specific QA datasets. Our experimental results demonstrate that the proposed approach achieves comparable or superior performance to previous approaches while improving data and computational efficiency. BALSa effectively mitigates audio hallucinations, further reinforcing its robustness in real-world applications. Beyond single-audio alignment, we extend to more complex audio understanding settings. In multi-audio scenarios, this extension further enhances modality alignment and leads to better performance on audio understanding and reasoning tasks. We further examine the role of discrimination-based training and observe that learning to handle multiple audio inputs alone is sufficient to improve overall understanding, without requiring explicit comparative objectives. Overall, this paper provides a scalable and efficient pathway for developing and advancing audio-aware large language models.